sammixboo (sammixboo):

Trying to help a certain someone. f(t) = -16t2 - 64t + 80 I know how to do this equation and I know the answer, just when this certain person and I try to solve it step by step we get a crazy answer

sammixboo (sammixboo):

f(t) = -16t2 - 64t + 80 0 = -16t2 - 64t + 80 0 = -16(t2 + 4t - 5) 0 = -16((t2 + 4t) - 5) So at this part, we either find the multiples of 5 that add up to 4, or we divide the constant of the t term by 2 then square our answer.

sammixboo (sammixboo):

If we divide the constant 4 by 2, we get 2. After we square the result we will get 4, so we add 4 inside the parenthesis and then subtract -16(4) from -5.. 0 = -16(t2 + 4t + 4) - 5 - (-16(4)))

sammixboo (sammixboo):

I feel like that is where I am getting confused, or is there factors of 5 that add up to 4??

jhonyy9 (jhonyy9):

t^2 +4t -5 =(t+5)(t-1)

sammixboo (sammixboo):

Oh, I get it, lol. 5 * -1

sammixboo (sammixboo):

5 + (-1) = 4, ok I don't need anymore help.

jhonyy9 (jhonyy9):

ok. was my pleasure good luck bye bye

sammixboo (sammixboo):

That took too much time to realize my mistake. I am way too tired Thanks